# How Do Developers Interact with AI? An Exploratory Study on Modeling Developer Programming Behavior

### 28 Mar 2026 v1

@@ 7.3. Implications for Practitioners
$\bullet$ Informing industrial AI assistant design: While we propose the S-IASE model in a research-based study, we expect this model can be applied for industry as well. We know the AI for coding companies, such as Cursor (Anysphere), and Codex (OpenAI) are collecting user behavior data to improve their models. Our proposed model may lead the way to systematically analyze user programming behavior and improve the models based on users’ real intention, action, and emotion.
@@
$\bullet$ Toward automated behavior annotation. Another promising direction is using our annotated dataset to train models to automatically detect parts of the S-IASE model, specifically the parts about intention, action, and support tools, and then prompt participants to reflect on their Emotions. This semi-automated pipeline could significantly reduce participant workload in future user studies, while enabling large-scale collection of fine-grained developer behavior data.
@@ 8. Related Work
8.1. AI Assistants Adoption In Software Engineering
The rapid adoption of AI assistants is transforming how developers perform software engineering tasks. Programming assistants like GitHub Copilot (GitHub, 2023b), GitHub Copilot Chat (GitHub, 2023a), ChatGPT (OpenAI, 2023a), Codex (OpenAI, 2023b), and Cursor (Cursor, 2023) are now widely integrated into software development environments. These AI assistants offer a variety of support from code comprehension, code suggestions, bug fixes, to documentation support.
Prior studies have explored the benefits, and challenges (Prather et al., 2023; Liang et al., 2024; Nam et al., 2024; Li et al., 2024) of AI assistants adoption in software engineering. For example, Liang et al. (Liang et al., 2024) found that developers primarily use AI assistants to reduce keystrokes, accelerate task completion, and recall syntax. They also reported that developers face challenges when AI-generated code fails to meet functional or non-functional requirements which leads low acceptance rates.
The influence between developers and AI assistants is mutual. Within the loop of software development process, developers not only actively make decisions on the top of AI outputs, but also their behavior is subtlety impacted by AI assistants in various activities, such as program comprehension and workflow management (Nguyen et al., 2024; Nam et al., 2024; Rajbhoj et al., 2024). This broader and also gradual impact suggests that AI assistants adoption in software engineering extends beyond individual programming tasks to shaping overall software development practices. In this work, we aim to expand this understanding by analyzing how developers’ intentions, actions, supporting tools and emotions interweave with AI assistance in practical software engineering tasks.
@@ 8.2. Developer-AI Interaction in Software Development
Previous studies have identified specific developer interaction patterns with AI assistants (Treude and Gerosa, 2025; Prather et al., 2023; Barke et al., 2023). For example, Barke et al. (Barke et al., 2023) adopted grounded theory to develop two modes of developer-AI interaction, one is in acceleration mode, the programmer knows what to do next and uses Copilot to get there faster; in exploration mode, the programmer is unsure how to proceed and uses Copilot to explore their options. Similarly, Prather et al. (Prather et al., 2023) also described two common patterns: shepherding (guiding AI suggestions) and drifting (exploring without clear goals). Amoozadeh et al. (Amoozadeh et al., 2024) not only found that CS1 students showed two similar patterns of ChatGPT use (using as a shortcut or iterative refinement), they also proposed a taxonomy of developer activities with ChatGPT, including reading, prompting, and verification.
However, these studies mainly derive interaction patterns from researchers’ external observations. In contrast, Mozannar et al. (Mozannar et al., 2024) directly examined developer–AI interaction from the developer’s perspective by asking participants to self-report their coding states. They developed a CUPS (CodeRec User Programming States) taxonomy with a specific focus on GitHub Copilot, such as deferring thought for later. They aim to categorize developer states during code completion with GitHub Copilot. Beyond physical or observed interaction (Treude and Gerosa, 2025), cognition such as trust perspective (Johnson et al., 2023; Sabouri et al., 2025; Wang et al., 2024) has also emerged as a key factor in developer-AI interaction. Choudhuri et al. (Choudhuri et al., 2024) examined how trust and system quality influenced developer interact with AI assistants like GitHub Copilot and ChatGPT. They found that functional value, ease of use, and perceived reliability significantly affected developers’ trust and willingness to adopt AI assistants. However, prior studies often treat interaction, behavior, and cognition separately, neglecting their dynamic interplay (Mozannar et al., 2024). Additionally, existing research typically constrains tool usage to specific AI assistants (Amoozadeh et al., 2024; Mozannar et al., 2024; Barke et al., 2023), limiting generalizability to real-world contexts.
Unlike the prior research, we permitted participants to select any AI assistants they commonly used, capturing developer-AI interaction patterns in their more natural environment. Our model describes developers’ intentions, subsequent actions with supporting tools, and emotion given a certain programming state. By this way, we aim to create a dynamic feedback loop that accurately reflects real-world software development.
@@ 9. Threats to Validity
Internal validity could be impacted by the splitting of participants into two groups. Participants were not randomly assigned, but were grouped based on their prior experience with AI. This introduces a potential selection bias. We also acknowledge the risk of aptitude bias, where observed behavioral patterns might be influenced by pre-existing participant traits or a higher baseline aptitude common among developers, rather than being solely dictated by AI usage. Thus, participants without AI experience may differ from those with AI experience in ways beyond just tool familiarity. Furthermore, although stratified sampling helped balance task-related expertise, it may also introduce residual bias due to differences in baseline programming skills, prior tool familiarity, or attitudes toward AI between groups. While we attempted to mitigate these risks through stratified sampling and post-hoc analysis, we acknowledge that residual confounding factors may remain. We used the pre-survey to help divide the participants into groups, ensuring that participants without or with little AI experience would not be using AI, and those with extensive experience would be using AI. Similarly, we used participant expertise to determine which activity they were assigned. Regarding the potential correlation between AI usage and overall aptitude with the tasks, participants were first explicitly assigned to the AI-assisted or non-AI groups based solely on their self-reported AI usage frequency. After this initial grouping, we employed stratified sampling based on participants’ self-reported programming expertise and task familiarity to assign them to one of the two tasks (PyT or JavaT). By doing so, we ensured that each task had participants with varying degrees of familiarity and expertise, thereby reducing the risk of aptitude bias. Nonetheless, we acknowledge that some residual effects from the correlation between AI usage frequency and overall task aptitude could remain as a potential limitation of our study design.
@@
Construct validity refers to whether we are asking the right questions in our user study procedure. To ensure the construct validity of the study, we carefully designed each stage of our procedure, including the survey questions, the activity tasks, and the interview questions. The questions and tasks were refined through extensive piloting (see Section 2.7). The self-annotation process relies on participants’ accurate recall and self-assessment. To minimize the impact on the results, we followed Mozannar’s practice (Mozannar et al., 2024) by instructing the participants to annotate their programming behavior right after the session. Participants were also allowed to create labels at their own discretion. There may be conceptual and technical differences between the two programming tasks (Java vs. Python). To mitigate, we collected participants’ task-related prior knowledge via the pre-survey and used it to assign tasks such that both tasks had varied distributions of expertise and familiarity. We acknowledge that Java and Python differ conceptually and technically, but the consistent patterns observed in both tasks suggest these differences did not meaningfully bias our results, so we analyzed them together. Moreover, we chose these tasks as they reflect real-world software development scenarios. Regarding emotion, we used a 7-point valence scale, following the practice of Gardella et al. (Gardella et al., 2024); however, we recognize that a 3-level score (Positive, Neutral, Negative) might reduce interpretation subjectivity. To mitigate the risk of the tasks could be directly solved by AI, we designed the tasks that require multi-step human involvement and environment-specific integration and we verified this via our pilot study. Thus we believe the threat is minimal. However, we acknowledge that the rapid evolution of LLMs to the solvability of these tasks may change over time. We acknowledge that the fixed 25-minute (determined through our pilot study) task may have introduced time pressure. However, we consider this a practical approximation of real-world development constraints. Our experiments showed this time-boxed window proved sufficient to capture a rich sequence of developer-AI interactions for our behavioral modeling.
@@
Conclusion validity is about the relationship between the treatments and the conclusions. Our work seeks to explore the potential of our model, develop a tool to help analyze developer behavior, and identify important behavioral patterns demonstrated by our participants. Our study was conducted under a fixed 25-minute time constraint, which may have amplified differences in action strategies between AI-assisted and non-AI groups. In particular, AI-assisted participants could more quickly generate and execute runnable code, whereas non-AI participants may have allocated relatively more time to reading and comprehension and searching the web for external resources. While our experiment setup reflects realistic time pressure in many programming scenarios, future work should investigate more action–tool relationships under no time constraints. We acknowledge that our validation and member-checking surveys reached a subset of participants, and we have analyzed the minority disagreements. We will explore and expand the explanations in future studies.
Finally, external validity refers to the generalizability of the research.
@@
One potential limitation is that participants were recruited primarily from a university setting rather than directly from industry. However, 75% of participants reported prior internship or industry experience, which mitigates this concern to some extent. We acknowledge that this population is better characterized as junior or early-career developers rather than experienced professionals, and thus it is not fully representative of expert developers in industry.
We believe many of the behaviors we observe are likely to generalize beyond students to early-career developers. Moreover, we expect our annotation tool to be adaptable for other coding tasks.
@@ 10. Conclusion and Future Work
In this work, we present a mixed-methods user study with 76 developers, performing programming tasks, to help develop deeper understanding of developer behavior when interacting with AI assistants. Through our user study, we propose a novel model for developer programming behavior that encapsulates four dimensions: Intention, Action, Supporting Tool, and Emotion. Our model provides insights that can be derived from snapshots of the data and also uncovered by a sequential view. We carefully designed and performed a series of user study experiments, including pilot study, pre-survey, programming activity with self-reporting, post-survey, and interview. Our findings include both aggregated and sequential behavioral patterns such as AI-assisted participants experiencing tension between productivity and comprehension.
Our study is not without limitations that open up opportunities for future work. Because our tasks emphasize completion within fixed 25-minute time, participants were not expected to produce production-ready, high-quality code. As a result, our observations capture the sequential patterns more than the end quality of artifacts. Future research should therefore investigate how the S-IASE model manifests in settings where code quality, maintainability, and long-term outcomes are central. Future research should also investigate these outliers to refine the granularity of the S-IASE model and further explore the socio-technical nuances of developer-AI interaction. Such work could clarify how AI assistance affects not only immediate developer experience but also the robustness and sustainability of software projects.
With all the valuable insights we uncovered, we advocate that future work conducts more experiments based on our proposed model to offer practical insights for integrating AI assistants into real-world software engineering processes.
@@
The data with personally identifiable information from participants that support the findings of this study are not publicly available due to privacy concerns and restrictions imposed by our Institutional Review Board (IRB). A replication package containing the package of our annotation tool, pre-survey, post-survey, the experiment scripts, and the anonymized results has been made available at: https://github.com/YinanWusoymilk/FSE-2026-How-Developers-Interact-with-AI.

### 26 Apr 2026 v2

@@ 7.3. Implications for Practitioners
$\bullet$ Informing industrial AI assistant design: While we propose the S-IASE model in a research-based study, we expect that this model can be applied in industry as well. We know the AI for coding companies continuously leverage user interaction data to refine their models. Our proposed model may lead the way to systematically analyze user programming behavior and improve the models based on users’ real intention, action, and emotion.
@@
$\bullet$ Toward automated behavior annotation. Another promising direction is using our annotated dataset to train models to automatically detect parts of the S-IASE model, specifically the parts about intention, action, and supporting tools, and then prompt participants to reflect on their Emotions. This semi-automated pipeline could significantly reduce participant workload in future user studies, while enabling large-scale collection of fine-grained developer behavior data.
@@ 8. Related Work
8.1. AI Assistant Adoption in Software Engineering
The rapid adoption of AI assistants is transforming how developers perform software engineering tasks. Programming assistants like GitHub Copilot (GitHub, 2023b), GitHub Copilot Chat (GitHub, 2023a), ChatGPT (OpenAI, 2023a), Codex (OpenAI, 2023b), and Cursor (Cursor, 2023) are now widely integrated into software development environments. These AI assistants offer a variety of support from code comprehension, code suggestions, bug fixes, to documentation support. Prior studies have explored the benefits, and challenges (Prather et al., 2023; Liang et al., 2024; Nam et al., 2024; Li et al., 2024) of AI assistant adoption in software engineering. For example, Liang et al. (Liang et al., 2024) found that developers primarily use AI assistants to reduce keystrokes, accelerate task completion, and recall syntax. They also reported that developers face challenges when AI-generated code fails to meet functional or non-functional requirements which leads to low acceptance rates.
The influence between developers and AI assistants is mutual. Within the loop of software development process, developers not only actively make decisions on top of AI outputs, but also their behavior is subtly impacted by AI assistants in various activities, such as program comprehension and workflow management (Nguyen et al., 2024; Nam et al., 2024; Rajbhoj et al., 2024). This broader and gradual impact suggests that AI assistant adoption in software engineering extends beyond individual programming tasks to shaping overall software development practices. In this work, we aim to expand this understanding by analyzing how developers’ intentions, actions, supporting tools, and emotions interweave with AI assistance in practical software engineering tasks.
@@ 8.2. Developer-AI Interaction in Software Development
Previous studies have identified specific developer interaction patterns with AI assistants (Treude and Gerosa, 2025; Prather et al., 2023; Barke et al., 2023). For example, Barke et al. (Barke et al., 2023) adopted grounded theory to develop two modes of developer-AI interaction, one is in acceleration mode, the programmer knows what to do next and uses Copilot to get there faster; in exploration mode, the programmer is unsure how to proceed and uses Copilot to explore their options. Similarly, Prather et al. (Prather et al., 2023) also described two common patterns: shepherding (guiding AI suggestions) and drifting (exploring without clear goals). Amoozadeh et al. (Amoozadeh et al., 2024) not only found that CS1 students showed two similar patterns of ChatGPT use (using as a shortcut or iterative refinement), they also proposed a taxonomy of developer activities with ChatGPT, including reading, prompting, and verification. However, these studies mainly derive interaction patterns from researchers’ external observations. In contrast, Mozannar et al. (Mozannar et al., 2024) directly examined developer–AI interaction from the developer’s perspective by asking participants to self-report their coding states. They developed a CUPS (CodeRec User Programming States) taxonomy with a specific focus on GitHub Copilot, such as deferring thought for later. They aim to categorize developer states during code completion with GitHub Copilot. Beyond physical or observed interaction (Treude and Gerosa, 2025), cognition such as trust perspective (Johnson et al., 2023; Sabouri et al., 2025; Wang et al., 2024) has also emerged as a key factor in developer-AI interaction. Choudhuri et al. (Choudhuri et al., 2025) examined how trust and system quality influenced developers’ interactions with AI assistants like GitHub Copilot and ChatGPT. They found that functional value, ease of use, and perceived reliability significantly affected developers’ trust and willingness to adopt AI assistants. However, prior studies often treat interaction, behavior, and cognition separately, neglecting their dynamic interplay (Mozannar et al., 2024). Additionally, existing research typically constrains tool usage to specific AI assistants (Amoozadeh et al., 2024; Mozannar et al., 2024; Barke et al., 2023), limiting generalizability to real-world contexts.
Unlike prior research, we permitted participants to select any AI assistants they commonly used, capturing developer-AI interaction patterns in their more natural environment. Our model describes developers’ intentions, subsequent actions with supporting tools, and emotions given a certain programming state. In this way, we aim to create a dynamic feedback loop that accurately reflects real-world software development.
@@ 9. Threats to Validity
Internal validity could be impacted because we split participants into two groups based on their prior experience with AI, and not randomly. This introduces a potential selection bias. We also acknowledge the risk of aptitude bias, where observed behavioral patterns might be influenced by pre-existing participant traits or a higher baseline aptitude common among developers, rather than being solely dictated by AI usage. Thus, participants without AI experience may differ from those with AI experience in ways beyond just tool familiarity. Furthermore, although stratified sampling helped balance task-related expertise, it may also introduce residual bias due to differences in baseline programming skills, prior tool familiarity, or attitudes toward AI between groups. While we attempted to mitigate these risks through stratified sampling and post-hoc analysis, we acknowledge that residual confounding factors may remain. We used the pre-survey to help divide the participants into groups, ensuring that participants without or with little AI experience would not be using AI, and those with extensive experience would be using AI. Similarly, we used participant expertise to determine which activity they were assigned. Regarding the potential correlation between AI usage and overall aptitude with the tasks, participants were first explicitly assigned to the AI-assisted or non-AI groups based solely on their self-reported AI usage frequency. After this initial grouping, we employed stratified sampling based on participants’ self-reported programming expertise and task familiarity to assign them to one of the two tasks (PyT or JavaT). By doing so, we ensured that each task had participants with varying degrees of familiarity and expertise, thereby reducing the risk of aptitude bias. Nonetheless, we acknowledge that some residual effects from the correlation between AI usage frequency and overall task aptitude could remain as a potential limitation of our study design.
@@
Construct validity refers to whether we are asking the right questions in our user study procedure. To ensure the construct validity of the study, we carefully designed each stage of our procedure, including the survey questions, the activity tasks, and the interview questions. The questions and tasks were refined through extensive piloting (see Section 2.7). The self-annotation process relies on participants’ accurate recall and self-assessment. To minimize the impact on the results, we followed Mozannar et al.’s practice (Mozannar et al., 2024) by instructing participants to annotate their programming behavior right after the session. Participants were also allowed to create labels at their own discretion. There may be conceptual and technical differences between the two programming tasks (Java vs. Python). To mitigate this, we collected participants’ task-related prior knowledge via the pre-survey and used it to assign tasks such that both tasks had varied distributions of expertise and familiarity. We acknowledge that Java and Python differ conceptually and technically, but the consistent patterns observed in both tasks suggest these differences did not meaningfully bias our results, so we analyzed them together. Moreover, we chose these tasks as they reflect real-world software development scenarios. Regarding emotion, we used a 7-point valence scale, following the practice of Gardella et al. (Gardella et al., 2024); however, we recognize that a 3-level score (Positive, Neutral, Negative) might reduce interpretation subjectivity. To mitigate the risk of the tasks that could be directly solved by AI, we designed the tasks that require multi-step human involvement and environment-specific integration and we verified this via our pilot study. Thus, we believe that the threat is minimal. However, we acknowledge that the rapid evolution of LLMs to the solvability of these tasks may change over time. We acknowledge that the fixed 25-minute task may have introduced time pressure. However, our pilot study determined this time-boxed window to be a practical approximation of real-world development constraints. Our experiments showed that this window was sufficient to capture a rich sequence of interactions for behavioral modeling.
@@
Conclusion validity is about the relationship between the treatments and the conclusions. Our work explores the potential of our model and identifies important behavioral patterns demonstrated by our participants. Our study was conducted with a 25-minute time constraint, which may have amplified differences in action strategies between AI-assisted and non-AI groups. In particular, AI-assisted participants more quickly generated and executed runnable code, whereas non-AI participants may have allocated more time to comprehension and searching. Future work should investigate more action–tool relationships without time constraints. We acknowledge that our validation and member-checking surveys reached a subset of participants, and we have analyzed the minority disagreements. We will explore and expand the explanations in future studies.
@@
External validity may be impacted by participant recruitment: we recruited primarily from a university setting, which may limit generalizability. However, 75% of participants reported prior internship or industry experience, which mitigates this concern to some extent. We acknowledge that this population is better characterized as junior or early-career developers rather than experienced professionals, and thus it is not fully representative of expert developers in industry. We believe that many of the behaviors we observe are likely to generalize beyond students to early-career developers. Moreover, we expect our annotation tool to be adaptable for other coding tasks.
@@ 10. Conclusion and Future Work
In this work, we present a mixed-methods user study with 76 developers, performing programming tasks, to help develop a deeper understanding of developer behavior when interacting with AI assistants. Through our user study, we propose a novel model for developer programming behavior that encapsulates four dimensions: Intention, Action, Supporting Tool, and Emotion. Our model provides insights that can be derived from snapshots of the data and uncovered by a sequential view. We carefully designed and performed a series of user study experiments, including pilot study, pre-survey, programming activity with self-reporting, post-survey, and interview. Our findings include both aggregated and sequential behavioral patterns such as AI-assisted participants experiencing tension between productivity and comprehension.
Our study opens up opportunities for future work. Because our tasks emphasized completion within a fixed time window, participants were not expected to produce production-ready code. Thus, our observations capture sequential patterns rather than artifact quality. Future research should investigate how the S-IASE model manifests in settings where code quality, maintainability, and long-term outcomes are central. Future research should also investigate these edge cases, such as the annotation ambiguities identified during member checking, to refine the granularity of the S-IASE model and further explore the socio-technical nuances of developer-AI interaction. Such work could clarify how AI assistance affects not only immediate developer experience but also the robustness and sustainability of software projects. With all the valuable findings we uncovered, we advocate that future work should conduct more experiments based on our proposed model to offer practical insights for integrating AI assistants into real-world software engineering processes.
@@
The data with personally identifiable information from participants that support the findings of this study are not publicly available due to privacy concerns and restrictions imposed by our Institutional Review Board (IRB). A replication package containing our annotation tool, pre-survey, post-survey, experiment scripts, and anonymized results is available at our GitHub repository: https://github.com/YinanWusoymilk/FSE-2026-How-Developers-Interact-with-AI.
Acknowledgments
This work is funded in part by NSF SHF #2006947 and #1749936.